Which is the equation with focus of (6,0)
A) Y=-1/6^2
B) y=1/12x^2
C) x=1/24y^2
D) x=-1/6y^2

Answer by lwsshak3(11628)

Which is the equation with focus of (6,0)
A) Y=-1/6^2
B) y=1/12x^2
C) x=1/24y^2
D) x=-1/6y^2
***
Parabola opens rightward:
Its basic equation: (y-k)^2=4p(x-h), (hk)=(x,y) coordinates of the vertex
For given equation:
vertex: (0,0)
axis of symmetry: y=0
p=6 (distance from vertex to focus on the axis of symmetry)
4p=24
equation: y^2=4px=24x
x=y^2/24 ans. C)